Find the area between the curves y^2=ax and ay^2=x^3(a>0), at the points of intersection other than the origin.
1
Expert's answer
2020-05-21T17:40:06-0400

Area between two curves f and g is given by:

A = "\int^b_a (f(y) -g(y)) dy" where a<y<b


Here,

g = y2=ax

"x= \frac{y^2}{a}"


f = ay2=x3

"x = a^{1\/3} y^{2\/3}"



A = "\int^a_{-a} (-\frac{y^2}{a} +a^{1/3} y^{2/3}) dy"


"=\int^a_{-a} -\frac{y^2}{a} dy + \int^a_{-a} a^{1/3} y^{2/3} dy"


"=\left. \frac{-y^3}{3a} \right|^{a}_{-a} + \left. \frac{3y^{5/3}a^{1/3}}{5} \right|^{a}_{-a}"


="-\frac{a^2}{3} + \frac{3a^2}{5} -\frac{a^2}{3} + \frac{3a^2}{5}"


= "-\frac{2a^2}{3} + \frac{6a^2}{5}"


= "\frac{-10a^2+18a^2}{15}"


="\frac{8a^2}{15}"